# Universal Transient Stability Analysis: A Large Language Model-Enabled Dynamics Prediction Framework

### v1

@@ 5 Conclusion
This paper introduces TSA-LLM, a universal framework that leverages a pre-trained LLM to achieve robust generalization across diverse OCs, unseen faults, and heterogeneous systems with a single architecture. This framework comprises a novel data processing pipeline featuring channel-independence, sample-wise normalization, and temporal patching to handle mixed-stability and long sequences from heterogeneous systems. Moreover, a parameter-efficient, freeze-and-finetune strategy for the augmented GPT architecture is designed to preserve generic sequence feature extraction from pre-trained T-blocks while allowing for TSA task adaptation. Furthermore, a two-stage fine-tuning scheme that combines TeaF with SchS is devised to mitigate cumulative prediction errors and ensuring long-horizon reliability. Case studies on two benchmark systems demonstrate that the proposed approach exhibits exceptional universality, low training and inference costs, and a considerable degree of interpretability, establishing it as a promising foundation for real-time TSA.
@@ Appendix A Generic Feature Extraction of T-Blocks
The decision to freeze the T-blocks in the GPT is justified by their optimization during pre-training, which molds them into powerful, generic feature extractors. Following the theoretical framework established in [zhou2023one, kim2021lipschitz], pre-training process is driven by the minimization of the attention gradient’s upper bound ($|G|_{2}$), as formulated in (26).
$\begin{aligned} |G|_{2}&\leq|A|_{2}\sum_{i}\left(a_{i,i}+\frac{1}{2}\right)\underbrace{\left|x_{i}-\sum_{j}a_{i,j}x_{j}\right|^{2}}_{\text{Self-alignment}}+\\ &|A|_{2}\sum_{i\neq j}a_{i,j}\underbrace{\left|x_{j}-\sum_{k}a_{i,k}x_{k}\right|^{2}}_{\text{Cross-alignment}}+\frac{|A|_{2}}{2}\sum_{i}\left|x_{i}\right|^{2}.\end{aligned}$ (26)
In (26), $A$ is the attention matrix, $x_{i}$, $x_{j}$ and $x_{k}$ are the token representation and $a_{i,j}$ signifies softmax probabilities from $A$. Minimizing this bound fosters two key alignment dynamics that function as an emergent form of dimensionality reduction. The Self-alignment term promotes feature stability by encouraging each feature vector $x_{i}$ to become more similar to its own attention-weighted context ($\sum a_{i,j}x_{j}$), which stabilizes the feature as they propagate through the network. Concurrently, the Cross-alignment term promotes feature co-direction by forcing feature vectors ($x_{i},x_{j}$) with high mutual attention to align with each other, mapping correlated inputs to representation vectors with similar directions. Together, these processes compel the model to represent complex sequences using a compact and consistent set of principal features. This inherent capability for generic and modality-agnostic feature extraction, forged during pre-training, is precisely why the T-block parameters are frozen for TSA tasks.

### v2

@@ 5 Conclusion
This paper introduces TSA-LLM, a universal framework that leverages a pre-trained LLM to achieve robust generalization across diverse OCs, unseen faults, and heterogeneous systems with a single architecture. This framework comprises a novel data processing pipeline featuring channel-independence, sample-wise normalization, and temporal patching to handle mixed-stability and long sequences from heterogeneous systems. Moreover, a parameter-efficient, freeze-and-finetune strategy for the augmented GPT architecture is designed to preserve generic sequence feature extraction from pre-trained T-blocks while allowing for TSA task adaptation. Furthermore, a two-stage fine-tuning scheme that combines TeaF with SchS is devised to mitigate cumulative pblackiction errors and ensuring long-horizon reliability. Case studies on two benchmark systems demonstrate that the proposed approach exhibits exceptional universality, low training and inference costs, and a considerable degree of interpretability, establishing it as a promising foundation for real-time TSA.
5.1 Scope and Limitations
The proposed framework is validated on several standard IEEE test systems with different scales and operating characteristics. Regarding system size, TSA-LLM has demonstrated transferability across multiple benchmark systems. Extending it to utility-scale grids will mainly depend on systematic simulations on larger and more heterogeneous systems to construct representative training and evaluation datasets. Regarding the diversity of operating conditions, the present experiments include randomized load variations and wind-induced stochasticity. Broader coverage will require additional renewable scenarios and more general non-optimal pre-fault states generated beyond OPF-centeblack settings. Regarding disturbance types, the current validation covers variation in fault location and contingency severity, but mainly under three-phase short-circuit faults. Extending the framework to a broader disturbance space will require incorporating additional symmetrical and asymmetrical faults, such as single line-to-ground and line-to-line faults. Regarding model assumptions, TSA-LLM follows a measurement-driven setting in which topology and parameter effects are learned implicitly from post-fault trajectories instead of being encoded explicitly. Relaxing this assumption will require introducing more explicit structural priors into the modeling framework.
5.2 Future Work
As future work, embedding specialized power system mechanism knowledge into the model structure or training objective (e.g., by adding physics informed regularization such as DAE/swing consistency residual losses) remains a promising direction worthy of further exploration, which may further enhance the framework’s interpretability and generalization under data-scarce or out-of-distribution scenarios. In addition, augmenting the input embedding with a topology-aware encoder (e.g., graph neural networks) to fuse explicit structural representations with sequence tokens is another promising extension. Identifying the optimal model scale for characterizing power system dynamics remains an open question and warrants further investigation.
@@ Appendix A Generic Feature Extraction of T-Blocks
The decision to freeze the T-blocks in GPT is grounded in their optimization during pre-training, which shapes them into generic sequence feature extractors. The self-attention input and mapping are first defined, followed by the gradient upper-bound result.
For a given channel index $i$ and sample index $j$, we follow the notation in the main text and denote the input to the T-block as the embedded token matrix $\mathbf{z}_{ij}^{0}$ defined in (9). To simplify the analysis below, we rewrite it as
$\mathbf{X}=\mathbf{z}_{ij}^{0}=(x_{1},\ldots,x_{P})^{\top}\in\mathbb{R}^{P\times d},$ (31)
where $P$ is the number of temporal patch tokens and, by a slight abuse of notation for notational simplicity, $x_{m}\in\mathbb{R}^{d}$ denotes the embedding of the $m$-th patch token (i.e., the $m$-th row of $\mathbf{z}_{ij}^{0}$). Define the self-attention mapping as
$f(\mathbf{X})=\mathrm{softmax}\!\left(\mathbf{X}A\mathbf{X}^{\top}\right)\mathbf{X},$ (32)
where $A=\mathbf{W}^{Q}(\mathbf{W}^{K})^{\top}\in\mathbb{R}^{d\times d}$. Let $\mathbf{P}=\mathrm{softmax}(\mathbf{X}A\mathbf{X}^{\top})$ with entries
$p_{m,n}=\frac{\exp(x_{m}^{\top}Ax_{n})}{\sum_{k=1}^{P}\exp(x_{m}^{\top}Ax_{k})}.$ (33)
Based on the gradient analysis of self-attention in [46, 14], the following bound holds.
Lemma A.1. Let $G=\left[\frac{\partial f(\mathbf{X})}{\partial x_{m}}\right]_{m=1}^{P}$ denote the Jacobian of $f(\mathbf{X})$ with respect to $\mathbf{X}$. Then,
$\begin{aligned} |G|_{2}\leq\;&|A|_{2}\sum_{m}\left(p_{m,m}+\frac{1}{2}\right)\underbrace{\left|x_{m}-\sum_{n}p_{m,n}x_{n}\right|^{2}}_{\text{Self-alignment}}\\ &+|A|_{2}\sum_{m\neq n}p_{m,n}\underbrace{\left|x_{n}-\sum_{k}p_{m,k}x_{k}\right|^{2}}_{\text{Cross-alignment}}+\frac{|A|_{2}}{2}\sum_{m}|x_{m}|^{2}.\end{aligned}$ (34)
Equation (34) provides a direct objective-level explanation for the effect of pre-training. During large-scale pre-training, parameter updates in self-attention tend to blackuce the gradient magnitude $|G|_{2}$; thus, optimization implicitly pushes down the right-hand side of (34). This induces two structublack alignment effects.
  • 1.
    First, the Self-alignment term $\left|x_{m}-\sum_{n}p_{m,n}x_{n}\right|^{2}$ shrinks the distance between each token representation $x_{m}$ and its attention-weighted context $\bar{x}_{m}=\sum_{n}p_{m,n}x_{n}$, which improves feature consistency after repeated T-block transformations.
  • 2.
    Second, the Cross-alignment term $\left|x_{n}-\sum_{k}p_{m,k}x_{k}\right|^{2}$ is weighted by $p_{m,n}$. Therefore, token pairs with larger attention weights are more strongly constrained to align with the same contextual center $\sum_{k}p_{m,k}x_{k}$, yielding co-directional representations for strongly correlated tokens.
Together, these alignment dynamics tend to project token representations onto a compact and consistent principal subspace. Since this mechanism operates on ordeblack token sequences without restricting the form of $X$, it is independent of semantic modality and can be regarded as modality-agnostic. For power system transient pblackiction, such structublack representation is particularly beneficial, as transient dynamics are typically dominated by a limited number of principal oscillatory modes. The alignment-induced projection concentrates trajectory information onto these dominant directions while suppressing blackundant components, thereby stabilizing long-range dependency modeling and mitigating error accumulation during iterative forecasting. This optimization-induced generic feature extraction behavior explains why freezing the pre-trained T-block parameters remains effective for TSA tasks.